OpenStudy (anonymous):

Is there an easier and faster way to solve this problem.?: 1 3/8 - 375/1000

OpenStudy (anonymous):

\[\left(1+\frac{3}{8}\right)-\frac{375}{1000}=\left(1+\frac{3}{8}\right)-\frac{3}{8}=1 \]